# Identifying Factual Inconsistencies in Summaries: Grounding Model Inference via Task Taxonomy

### v2

@@ 5.3 Supervised Analysis
Fine-Grained Evaluation
Error Type Predictions
@@ 6 Conclusion
7 Ethical Statements
This work does not pose direct ethical concerns, as we utilize existing datasets on a well-studied task. However, we do emphasize that factual predictions by our proposed approaches should not be taken without verification. We provide failed examples by zero-shot LLM inference in Appx. F.
@@ 7 Ethical Statements
Second, both the zero-shot paradigm and supervised paradigm may not fully capture the nuances of complex summaries. We list concrete qualitative examples in Appendix F on the failed cases by LLMs. Specifically for zero-shot paradigm, the failed cases could come from imperfect instruction following, as well as ambiguous descriptions of the task taxonomy that are not fully comprehensive. For the supervised paradigm, it indeed requires either human annotated examples, or synthetic data generation, which may not generalize as well as the zero-shot inference.
@@ Appendix C Zero-Shot Experimental Settings
For FacTax-WD that operates on summary windows, $n$ important sentences are extracted independently to maximize the combined recall of ROUGE-1 and ROUGE-2 metrics in relation to the summary Table 8 shows the alignment thresholds we adopted for the two datasets.
Summ-Alignment Window-Alignment ($n$=5)
@@
Table 8: The maximum length of aligned context for FacTax, and the averaged length of aligned context per summary window for FacTax-WD. $n$ is the number of sentences extracted for each summary window. For SQuALITY, some of the retrieved sentences can be quite short.
@@
Five non-LLM baselines, QuestEval, QaFactEval, SummaC-ZS, AlignScore and Align produce a consistency score for each summary, which requires a threshold to convert to the classification label. For each dataset and origin split in AggreFact-FtSota test set, following Tang et al. (2023), we tune the threshold to reach the best balanced accuracy on the corresponding part of the dataset or origin split in AggreFact-FtSota validation set.
For DiaSumFact, we use validation split of AggreFact-FtSota dataset to tune the threshold for the baselines.
@@ Appendix D Supervised Experimental Settings
Appendix E AI Assisted Writing
ChatGPT from OpenAI is used in this writing for revising certain expressions.
Appendix F Failure Cases by LLMs

### v3

@@ 5.3 Supervised Analysis
Fine-Grained Evaluation
Error Type Predictions
@@ 6 Conclusion
Second, both the zero-shot paradigm and supervised paradigm may not fully capture the nuances of complex summaries. We list concrete qualitative examples in Appendix E on the failed cases by LLMs. Specifically for zero-shot paradigm, the failed cases could come from imperfect instruction following, as well as ambiguous descriptions of the task taxonomy that are not fully comprehensive. For the supervised paradigm, it indeed requires either human annotated examples, or synthetic data generation, which may not generalize as well as the zero-shot inference.
Acknowledgments
We are grateful to the reviewers for their insightful comments. Liyan Xu was supported in part by the National Science Foundation grant IIS-2303678.
@@ Appendix C Zero-Shot Experimental Settings
For FacTax-WD that operates on summary windows, $n$ important sentences are extracted independently to maximize the combined recall of ROUGE-1 and ROUGE-2 metrics in relation to the summary. Table 8 shows the alignment thresholds we adopted for the two datasets.
For non-LLM systems, the context length limit is usually shorter (512 for BERT models). We perform window-based inference accordingly, so that the aligned context for each summary window falls within 512 tokens. The final result is thereby aggregated through all windows.
FacTax FacTax-WD ($n$=5)
@@
Table 8: The maximum length of aligned context for FacTax, and the averaged length of aligned context per summary window for FacTax-WD, with $n$ being the number of sentences extracted for each summary window. For SQuALITY, some of the retrieved sentences can be quite short.
@@
Five non-LLM baselines, QuestEval, QaFactEval, SummaC-ZS, AlignScore and Align produce a consistency score for each summary, which requires a threshold to convert to the classification label. For each dataset in the AggreFact-FtSota test set, following Tang et al. (2023), we tune the threshold to reach the best balanced accuracy on the corresponding AggreFact-FtSota validation set.
For DiaSumFact without a specific validation set, we use the AggreFact-FtSota validation set to tune the threshold for the non-LLM baselines.
@@ Appendix D Supervised Experimental Settings
Appendix E Failure Cases by LLMs